write an equation of the line of 9, -9 and parralle to the graph 5x- 9 y =4
Answer by mananth(16946)

5 x -9 y = 4
Find the slope of this line
make y the subject
-9y=-5x+4
Divide by -9
y = (5/ 9)x- 4/9
Compare this equation with y=mx+b
slope m = 5/ 9
The slope of a line parallel to the above line will be the same
The slope of the required line will be 5/ 9
m= 5/ 9 ,point ( 9 , -8 )
Find b by plugging the values of m & the point in
y=mx+b
-8 = 5 + b
b= -14
m= 5/9
Plug value of the slope and b in y = mx +b
The required equation is y = 5/9 x-14

m.ananth@hotmail.ca